### v2

# TIP-Editor: An Accurate 3D Editor Following Both Text-Prompts And Image-Prompts

Abstract.
Text-driven 3D scene editing has gained significant attention owing to its convenience and user-friendliness. However, existing methods still lack accurate control of the specified appearance and location of the editing result due to the inherent limitations of the text description. To this end, we propose a 3D scene editing framework, TIP-Editor, that accepts both text and image prompts and a 3D bounding box to specify the editing region. With the image prompt, users can conveniently specify the detailed appearance/style of the target content in complement to the text description, enabling accurate control of the appearance. Specifically, TIP-Editor employs a stepwise 2D personalization strategy to better learn the representation of the existing scene and the reference image, in which a localization loss is proposed to encourage correct object placement as specified by the bounding box. Additionally, TIP-Editor utilizes explicit and flexible 3D Gaussian splatting (GS) as the 3D representation to facilitate local editing while keeping the background unchanged. Extensive experiments have demonstrated that TIP-Editor conducts accurate editing following the text and image prompts in the specified bounding box region, consistently outperforming the baselines in editing quality, and the alignment to the prompts, qualitatively and quantitatively.
submissionid: 820copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; July 28–August 01, 2024; Denverisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Rendering
Figure 1. TIP-Editor excels in precise and high-quality localized editing given a 3D bounding box, and allows the users to perform various types of editing on a 3D scene, such as object insertion $(a,b)$, whole object replacement $(d)$, part-level object editing $(e)$, combination of these editing types (i.e. sequential editing, $(c)$), and stylization $(f)$. The editing process is guided by not only the text but also one reference image, which serves as the complement of the textual description and results in more accurate editing control. Images in the text prompts denote their associated rare tokens, which are fixed without optimization.
1. Introduction
Due to the unprecedented photorealistic rendering quality, methods that use radiance field-related representations (e.g. NeRF (Mildenhall et al., 2021) and 3D Gaussian Splatting (Kerbl et al., 2023)) have been more and more popular in 3D reconstruction field (Yu et al., 2021; Li et al., 2021; Cao, 2023) and various downstream 3D editing tasks, such as texture editing (Xiang et al., 2021; Richardson et al., 2023), shape deformation (Yang et al., 2022; Xu and Harada, 2022), scene decomposition (Tang et al., 2022), and stylization (Wang et al., 2023b).
Generative editing, which only requires high-level instructions (e.g. text prompts), emerges as a new approach in complement to previous painting-like and sculpting-like editing approaches (Xiang et al., 2021; Yang et al., 2022) that require extensive user interactions. Among these methods, text-driven methods (Haque et al., 2023; Zhuang et al., 2023) have gained significant attention due to their convenience and have achieved remarkable progress due to the success of large-scale text-to-image (T2I) models.
However, methods using only text as the condition struggle to precisely generate editing results with the specified appearance at the specified location due to the inherent limitations of the text description. For example, existing text-driven methods usually produce less satisfactory results (Fig. 6) if we want to dress the toy in a special heart-shaped sunglasses or give the male the Joker makeup appeared in the movie The Dark Knight. Moreover, it is hard to specify the accurate editing location by text guidance (Fig. 7). These challenges primarily stem from the diverse appearances of the generated objects and the diverse spatial layout of the generated scenes.
To overcome the challenges above, we present TIP-Editor, which allows the users to intuitively, conveniently, and accurately edit the exiting GS-based (3D Gaussian Splatting (Kerbl et al., 2023)) radiance fields using both text prompts and image prompts. Our framework achieves such capabilities through two crucial designs. (1) The first one is a novel stepwise 2D personalization strategy that enables precise appearance control (via a reference image) and location control (via a 3D bounding box). Specifically, it contains a scene personalization step, which includes a localization loss to ensure the editing occurs inside the user-defined editing region, and a separate novel content personalization step dedicated to the reference image based on LoRA (Hu et al., 2021). (2) The second one is adopting explicit and flexible 3D Gaussian splatting (Kerbl et al., 2023) as the 3D representation since it is efficient and, more importantly, highly suitable for local editing.
We conduct comprehensive evaluations of TIP-Editor across various real-world scenes, including objects, human faces, and outdoor scenes. Our editing results (Fig. 1 and Fig. 3) successfully capture the unique characteristics specified in the reference images. This significantly enhances the controllability of the editing process, presenting considerable practical value. In both qualitative and quantitative comparisons, TIP-Editor consistently demonstrates superior performance in editing quality, visual fidelity, and user satisfaction when compared to existing methods.
Our contributions can be summarized as follows:
  • We present TIP-Editor, a versatile 3D scene editing framework that allows the users to perform various editing operations (e.g. object insertion, object replacement, re-texturing, and stylization) guided by not only the text prompt but also by a reference image.
  • We present a novel stepwise 2D personalization strategy, which features a localization loss in the scene personalization step and a separate novel content personalization step dedicated to the reference image based on LoRA, to enable accurate location and appearance control.
  • We adopt 3D Gaussian splatting to represent scenes due to its rendering efficiency and, more importantly, its explicit point data structure, which is very suitable for precise local editing.
2. Related Works
Text-guided image generation and editing
Text-to-image (T2I) diffusion models  (Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022a), trained on large-scale paired image-text datasets, have gained significant attention since they can generate diverse and high-quality images that match the complicated text prompt. Instead of directly generating images from scratch, another popular and closely related task is to edit the given image according to the text prompt (Meng et al., 2021; Couairon et al., 2022; Kawar et al., 2022; Hertz et al., 2022; Avrahami et al., 2022; Brooks et al., 2022).
Another popular task is object/concept personalization, which aims at generating images for a specified object/concept defined in the given image collection. Textual Inversion (TI) (Gal et al., 2022a) optimizes special text token(s) in the text embedding space to represent the specified concept. DreamBooth (Ruiz et al., 2022) fine-tunes the entire diffusion model with a class-specific prior preservation loss as regularization. In general, DreamBooth generates higher-quality images since it involves a larger amount of updated model parameters (i.e. the whole UNet model). However, all the aforementioned methods do not support generating images containing multiple personalized objects simultaneously.
Custom Diffusion (Kumari et al., 2023) extends the above task to generate multiple personalized concepts in one image simultaneously. Although separate special text tokens are assigned to each concept, the UNet is updated by all concepts, resulting in less satisfactory personalization results. Furthermore, it lacks a localization mechanism to specify the interaction between two concepts (Fig. 8). In contrast, we propose a stepwise 2D personalization strategy to learn the existing scene and the new content separately, achieving high-quality and faithful personalization results and being generalizable to sequential editing scenarios.
Radiance field-based 3D object/scene generation
The success of T2I diffusion models has largely advanced the development of 3D object/scene generation. One seminal contribution, DreamFusion (Poole et al., 2022), introduces score distillation sampling (SDS), which distills knowledge from a pre-trained 2D T2I model to optimize a radiance field without the reliance on any 3D data. Most of the subsequent works adopt such an optimization-based pipeline and make further progresses by introducing an extra refinement stage (e.g., Magic3D (Lin et al., 2022) and DreamBooth3D (Raj et al., 2023)), or proposing more suitable SDS variants (e.g., VSD (Wang et al., 2023c)), or using more powerful 3D representations (Chen et al., 2023b; Yi et al., 2023; Chen et al., 2023d), such as GS (Kerbl et al., 2023).
Furthermore, a body of research (Deng et al., 2022; Melas-Kyriazi et al., 2023; Tang et al., 2023; Qian et al., 2023) endeavors to integrate reference images within the optimization framework. This integration is facilitated by various techniques, including the application of reconstruction loss, employment of predicted depth maps, and the execution of a fine-tuning process. Nevertheless, these methods are constrained to generate a single object from scratch and cannot edit existing 3D scenes.
Radiance field-based 3D object/scene editing
Earlier works (Wang et al., 2022, 2023b) mainly focus on global style transformation of a given 3D scene, which takes text prompts or reference images as input and usually leverage a CLIP-based similarity measure (Radford et al., 2021) during optimization. Several studies enable local editing on generic scenes by utilizing 2D image manipulation techniques (e.g. inpainting) (Liu et al., 2022; Kobayashi et al., 2022; Bao et al., 2023) to obtain new training images to update the existing radiance field. Some works adopt 3D modeling techniques (e.g. mesh deformation (Yuan et al., 2022; Yang et al., 2022; Xu and Harada, 2022; Wang et al., 2023a) or point clouds deformation  (Wu et al., 2023; Chen et al., 2023c)) to propagate the shape deformation to the underlying radiance field. However, these methods require extensive user interactions.
Recently, text-driven radiance field editing methods have gained more and more attention for their editing flexibility and accessibility. For example, Instruct-NeRF2NeRF (Haque et al., 2023) employs an image-based diffusion model (InstructPix2Pix (Brooks et al., 2022)) to modify the rendered image by the users’ instructions, and subsequently update the 3D radiance field with the modified image. DreamEditor (Zhuang et al., 2023), 3D Paintbrush (Decatur et al., 2023) and Vox-E (Sella et al., 2023) enable better local editing by adopting explicit 3D representations (i.e. mesh and voxel), where the editing region is automatically determined by the 2D cross-attention maps. GaussianEditor (Chen et al., 2023a; Fang et al., 2023) adopts GS as the scene representation and incorporates 3D semantic segmentation (Kirillov et al., 2023a; Cen et al., 2023) to facilitate efficient and precise scene editing. However, these text-driven approaches lack precise control over the specified appearance and position of the editing results.
A concurrent work, CustomNeRF (He et al., 2023), is most related to our task setting. But CustomNeRF only supports the object replacement task, since it requires an object that can be detected by the segmentation tool (Kirillov et al., 2023b) existing in the implicit NeRF scene, as the editing target. In contrast, we adopt explicit GS as the 3D representation which facilitates our method to perform more editing tasks (e.g., object insertion and stylization).
3. Background
3D Gaussian Splatting.
3D Gaussian Splatting (GS) (Kerbl et al., 2023) quickly draws tremendous attention due to its high rendering quality and efficiency. GS utilizes a set of point-like anisotropic Gaussians $g_{i}$ to represent the scene: $\mathcal{G}=\{g_{1},g_{2},...,g_{N}\}$. Each $g_{i}$ contains a series of optimizable attributes, including center position $\mu\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}^{1}$, 3D covariance matrix $\Sigma$, and color $c$. The differentiable splatting rendering process is outlined as follows:
(1) $C=\sum_{i\in\mathcal{N}}c_{i}\sigma_{i}\prod_{i-1}^{j=1}(1-\sigma_{j}),\quad\sigma_{i}=\alpha_{i}G(x)=\alpha_{i}e^{-\frac{1}{2}(x)^{T}\Sigma^{-1}(x)},$
where $j$ indexes the Gaussians in front of $g_{i}$ according to their distances to the optical center in ascending order, $\mathcal{N}$ is the number of Gaussians that have contributed to the ray, and $c_{i}$, $\alpha_{i}$, and $x_{i}$ represent the color, density, and distance to the center point of the $i$-th Gaussian, respectively.
Optimizing Radiance Fields with SDS Loss.
Score distillation sampling (SDS)  (Poole et al., 2022) optimizes a radiance field by distilling the priors from a Text-to-Image (T2I) diffusion model for 3D generation. The pre-trained diffusion model $\phi$ is used to predict the added noise given a noised image $\hat{I}_{t}$ and its text condition $y$.
(2) $\nabla_{\theta}\mathcal{L}_{SDS}(\phi,\hat{I}=f(\theta))=\mathbb{E}_{\epsilon,t}\bigg{[}w(t)(\epsilon_{\phi}(\hat{I}_{t};y,t)-\epsilon)\frac{\partial\hat{I}}{\partial\theta}\bigg{]},$
where $\theta$ denotes the parameters of the radiance field, $f(\cdot)$ is the differentiable image formation process, and $w(t)$ is a predefined weighting function derived from noise level $t$.
Figure 2. Method overview. TIP-Editor optimizes a 3D scene that is represented as 3D Gaussian splatting (GS) to conform with a given hybrid text-image prompt. The editing process includes three stages: 1) a stepwise 2D personalization strategy, which features a localization loss in the scene personalization step and a separate novel content personalization step dedicated to the reference image based on LoRA (Sec. 4.1); 2) a coarse editing stage using SDS (Sec. 4.2); and 3) a pixel-level texture refinement stage, utilizing carefully generated pseudo-GT image from both the rendered image $I_{c}$ and the denoised image $I_{c}^{d}$ (Sec. 4.3).
4. Method
Given posed images (i.e., images and their associated camera parameters estimated by COLMAP (Schönberger and Frahm, 2016)) of the target scene, our goal is to enable more accurate editing following a hybrid text-image prompt within a user-specified 3D bounding box. We choose 3D Gaussian splatting (GS) (Kerbl et al., 2023) to represent the 3D scene since GS is an explicit and highly flexible 3D representation method, which is beneficial for the following editing operations, especially local editing.
As shown in Fig. 2, TIP-Editor contains three major steps, including 1) a stepwise 2D personalization of the existing scene and the novel content (Sec. 4.1), 2) a coarse 3D editing stage using score distillation sampling (SDS) (Poole et al., 2022) (Sec. 4.2), and 3) a pixel-level refinement of the 3D scene (Sec. 4.3).
4.1. Stepwise 2D Personalization
In general, our stepwise personalization of the pre-trained T2I model (i.e., Stable Diffusion (SD) (Rombach et al., 2022a)) is based on DreamBooth (Ruiz et al., 2022), but with two significant modifications. These changes are essential to personalize both the existing scene and the novel content in the reference image. First, in the 2D personalization of the existing scene, we propose an attention-based localization loss to enforce the interaction between the existing and the novel content specified by the provided 3D bounding box (e.g., sunglasses on the forehead, see Fig. 7). Note that the reference image is not involved in this step. Second, in the 2D personalization of the novel content, we introduce LoRA layers to better capture the unique characteristics of the specified item in the reference image.
2D personalization of the existing scene.
We first personalize the SD to the given scene to facilitate various types of editing of the scene afterward. Specifically, the initial text prompt (e.g. ”a toy”) is obtained using an image captioning model, BLIP-2 (Li et al., 2023). To enhance the specificity of the scene, we add a special token $V_{1}$ in front of the noun describing the scene, resulting in a scene-specific text prompt (e.g., “a $V_{1}$ toy”) as in (Zhuang et al., 2023). The UNet $\epsilon_{\phi}$ of the T2I model is fine-tuned with the reconstruction loss and the prior preservation loss (Ruiz et al., 2022). The input of the reconstruction training includes the scene-specific text and a rendered image of the 3D scene from a random view. The input of the prior preservation training includes the initial text and a random image generated by SD using the initial text as input (omitted in Fig. 2 to reduce clutter). The above losses are computed as follows:
(3) $\begin{split}\mathcal{L}_{scene}=&\mathbb{E}_{z,y,\epsilon,t}||\epsilon_{\phi_{1}}(z_{t},t,p,y)-\epsilon||_{2}^{2}+\\ &\mathbb{E}_{z^{*},y^{*},\epsilon,t^{*}}||\epsilon_{\phi_{1}}(z^{*}_{t},t^{*},p^{*},y^{*})-\epsilon||_{2}^{2}\end{split}$
where $y$ denotes the text prompt, $t$ the timestep, $z_{t}$ the noised latent code at $t$-th timestep extracted from the input scene image, and $p$ the camera pose. Superscript * denotes the corresponding variables used in prior preservation training. Note that we add an additional camera pose $p$ to the condition embeddings in the network to have a better viewpoint control of the generated images from the SD, facilitating the subsequent SDS-based 3D scene optimization. Since randomly generated images for prior preservation training do not have a meaningful “scene pose”, we assign a fixed camera pose $p^{*}=I_{4}$ that will never be used for rendering.
To encourage accurate localization of the target object, we introduce an attention-based localization loss (Fig. 2) during the 2D scene personalization to encourage the SD to generate images containing the required scene-object interaction. This step is particularly important if the target object is specified at a rarely seen location (e.g., sunglasses on the forehead, see Fig. 7). The actual location of the target object generated by SD is extracted from the cross-attention map $A_{t}$ of the object keyword (e.g., “sunglasses”) following (Hertz et al., 2022). The wanted location of the target object (i.e., GT editing region) is obtained by projecting the provided 3D bounding box to the image plane. The loss between the actual and the wanted location is defined as:
(4) $\mathcal{L}_{loc}=(1-\underset{s\in\mathcal{S}}{max}(A_{t}^{s}))+\lambda\sum_{s\in\bar{\mathcal{S}}}||A_{t}^{s}||_{2}^{2}$
where, $\lambda$ is a weight to balance two terms, $\mathcal{S}$ the GT editing mask region (projection of the 3D bounding box $\mathcal{B}$) and $\bar{\mathcal{S}}$ the otherwise. Intuitively, this loss encourages a high probability inside the editing area and penalizes the presence of the target object outside the editing area. As demonstrated in our ablation study (Fig. 7), this loss is crucial for ensuring precise editing within the specified region.
2D personalization of the novel content.
We introduce a dedicated personalization step using LoRA (Hu et al., 2021) (with the UNet fixed) to better capture the unique characteristics contained in the reference image. This step is essential to reduce the negative influence (e.g. concept forgetting (Kumari et al., 2023)) when learning (personalizing) multiple concepts, resulting in a better representation of both the scene and the novel content. Specifically, we train the additional LoRA layers inserted to the previously personalized and fixed T2I model $\epsilon_{\phi^{*}}$. Similar to the last step, we obtain the initial text prompt using BLIP-2 model and insert a special token $V_{2}$ into it, yielding an object-specific text prompt $y^{r}$ of the reference object (e.g. “$V_{2}$ sunglasses”). The new LoRA layers are trained with the following loss function:
(5) $\mathcal{L}_{ref}=\mathbb{E}_{z^{r},y^{r},\epsilon,t}||\epsilon_{\phi_{2}}(z_{t}^{r},t,p^{*},y^{r})-\epsilon||_{2}^{2}$
After training, the content of the scene and the reference image are stored in UNet and added LoRA layers, respectively, resulting in largely reduced mutual interference.
4.2. Coarse Editing via SDS Loss
We optimize the selected Gaussians $\mathcal{G^{\mathcal{B}}}\in\mathcal{B}$ (i.e., those inside the bounding box $\mathcal{B}$) with SDS loss from the personalized T2I diffusion model $\epsilon_{\phi_{2}}$. Specifically, we input randomly rendered images $\hat{I}$ using sampled camera poses $p$ and the text prompt $y^{G}$ into the T2I model $\epsilon_{\phi_{2}}$, and calculate the global scene SDS Loss as follows:
(6) $\nabla_{\mathcal{G}}\mathcal{L}_{SDS}^{G}(\phi_{2},f(\mathcal{G}))=\mathbb{E}_{\epsilon,t}\bigg{[}w(t)(\epsilon_{\phi_{2}}(z_{t};t,p,y^{G})-\epsilon)\frac{\partial z}{\overset{}{\partial}\hat{I}}\frac{\partial\hat{I}}{\overset{}{\partial}\mathcal{G}}\bigg{]}$
where $y^{G}$ is the text prompt including special tokens $V_{1},V_{2}$ and describes our wanted result, $f(\cdot)$ the GS rendering algorithm.
It is noteworthy that the selection and update criteria of the Gaussians $\mathcal{G^{\mathcal{B}}}$ to be optimized are slightly different for different types of editing tasks. For object insertion, we duplicate all the Gaussians inside the bounding box and exclusively optimize all the attributes of these new Gaussians. For object replacement and re-texturing, all the Gaussians inside the bounding box will be updated. For stylization, optimization is applied to all the Gaussians in the scene. Note that we only update the colors (i.e., the spherical harmonic coefficients) for re-texturing instead of updating all the attributes.
Since the foreground and background of a GS-based scene are readily separable given the bounding box $\mathcal{G^{\mathcal{B}}}$, we introduce another local SDS loss for object-centric editing (e.g., object insertion/replacement) to reduce artifacts as follows:
(7) $\nabla_{\mathcal{G^{\mathcal{B}}}}\mathcal{L}_{SDS}^{L}(\phi_{2},f(\mathcal{G^{\mathcal{B}}}))=\mathbb{E}_{\epsilon,t}\bigg{[}w(t)(\epsilon_{\phi_{2}}(z_{t};t,p,y^{L})-\epsilon)\small{\frac{\partial z}{\overset{}{\partial}\hat{I}}\frac{\partial\hat{I}}{\overset{}{\partial}\mathcal{G^{\mathcal{B}}}}}\bigg{]}$
where $y^{L}$ is the text prompt including the special tokens $V_{2}$ and only describes our wanted new object, $\hat{I}$ the rendered images containing only the foreground object.
We employ $\mathcal{L}_{SDS}^{G}$ and $\mathcal{L}_{SDS}^{L}$ with weight $\gamma$ to optimize $\mathcal{G^{\mathcal{B}}}$:
(8) $\mathcal{L_{\mathcal{SDS}}}=\gamma\mathcal{L}_{SDS}^{G}+(1-\gamma)\mathcal{L}_{SDS}^{L}$
4.3. Pixel-Level Image Refinement
In this stage, we introduce a pixel-level reconstruction loss to effectively enhance the quality of the editing results, since the 3D results directly optimized with SDS loss usually contain artifacts (e.g. green noise on the glasses’ frame, needle-like noise on the hair in Fig. 11).
The core of this stage is to create a pseudo-GT image $I_{gt}$ to supervise the rendered image $I_{c}$ from the coarse GS. Firstly, we follow SDEdit (Meng et al., 2021) to add noise on the rendered image $I_{c}$ to obtain the noised image $I_{c}^{d}$ and then utilized the personalized T2I model $\epsilon_{\phi_{2}}$ as a denoising network and obtain the denoised image $I_{c}^{d}$. The denoising process effectively reduces the artifacts in $I_{c}$ (see Fig.  D.1 in the supplementary), but also alters the background image. Secondly, we obtain the binary instance mask $M^{inst}$ of the edited object/part by rendering only the editable Gaussians $\mathcal{G}^{\mathcal{B}}$ and thresholding its opacity mask. Then, we render a background image $I_{bg}$ with only the fixed Gaussians. Finally, the pseudo-GT image $I_{gt}$ is obtained as:
(9) $I_{gt}=M^{inst}\odot I_{c}^{d}+(1-M^{inst})\odot I_{bg}$
This process ensures that the background image is clean and the same as the original scene while the foreground editable region is enhanced by the T2I model $\epsilon_{\phi_{2}}$. Using this pseudo-GT image as pixel-level supervision effectively enhances the resultant texture and reduces floaters (Fig. 11). MSE loss is applied between the rendered image $I_{c}$ and the created pseudo-GT image $I_{gt}$. A flowchart depicting the complete preparation of $I_{gt}$ is included in the supplementary.
To maintain better coverage, the rendering camera poses cover all elevation and azimuth angles with an interval of $30^{\circ}$ within a predefined range. To maintain better view-consistency of the denoised images, we set a small noise level ($t_{0}=0.05$, .i.e, “intermediate time” in SDEdit). Using such a small noise level effectively enhances fine texture details, removes small artifacts, and does not introduce significant shape and appearance change, maintaining better view consistency for the target editing region.
5. Experiments
Figure 3. Editing results of the proposed TIP-Editor. Images in the text prompts denote their associated rare tokens, which are fixed without optimization.
5.1. Experimental Setup
Implementation Details
We use the official code to train the original scene GS, with the default hyper-parameters. In the stepwise 2D personalization stage, the scene personalization step consists of 1k iterations, while the novel content personalization contains 500. We set $\lambda=0.1$ in $\mathcal{L}_{loc}$. In the coarse editing stage, we adopt the sampling strategy of views from  (Zhuang et al., 2023). The size of the rendered images is 512$\times$512. Owing to the different complexity of the editing task, this stage requires optimizing for 1Ksimilar-$\sim$5K iterations, consuming approximately 5$\sim$25 minutes. The refinement stage takes 3K iterations with the supervision of the generated $I_{gt}$, concluding in less than 3 minutes. More implementation details can be found in the supplementary.
Dataset
To comprehensively evaluate our method, We select six representative scenes with different levels of complexity following previous works (Zhuang et al., 2023; Haque et al., 2023; Chen et al., 2023a). These scenes include objects in simple backgrounds, human faces, and complex outdoor scenes. For each editing, a hybrid prompt, consisting of text and a reference image obtained from the Internet, is employed to guide the editing. Additionally, we manually set a 3D bounding box to define the editing region.
Baselines
Due to the lack of dedicated image-based editing baselines, we compare with two state-of-the-art text-based radiance field editing methods, including Instruct-NeRF2NeRF (“I-N2N”) (Haque et al., 2023) and DreamEditor (Zhuang et al., 2023). I-N2N utilizes Instruct-pix2pix (Brooks et al., 2022) to update the rendered multi-view images according to special text instructions. DreamEditor adopts a mesh-based representation and includes an attention-based localization operation to support local editing. For a fair comparison, we replace its automatic localization with a more accurate manual selection. See our supplementary for more implementation details.
Figure 4. Sequential editing results. We show two rendered images of the 3D scene after every editing step, indicated by the number in the top-left corner. $V_{*}$, $V_{**}$, and $V_{***}$ represent the special tokens of the scene in different sequences of editing.
Figure 5. Results of using a generated image as the reference. We first generate several candidate images by the diffusion model using text prompts, then we choose one as the reference image for editing.
Evaluation Criteria
For quantitative evaluation, we adopt CLIP Text-Image directional similarity following (Haque et al., 2023; Zhuang et al., 2023) to assess the alignment of the editing outcomes with the given text prompt. To evaluate image-image alignment (between the edited scene and the reference image), we follow (He et al., 2023) to calculate the average DINO similarity (Oquab et al., 2023) between the reference image and the rendered multi-view images of the edited 3D scene. Detailed information about these calculations is available in the supplementary.
Additionally, we conduct a user study and ask the participants (50 in total) to evaluate the results of different methods from two aspects (overall “Quality”, and “Alignment” to the reference image). The user study includes 10 questions, each containing the edited results of the two baselines and ours rendered into rotating videos in random order (see our supplementary). The 10 questions have covered various scenes and editing types to better compare the methods under different scenarios.
5.2. Visual Results of TIP-Editor
In Fig.1 and Fig. 3, we present qualitative results of TIP-Editor. Video demonstrations are included in the supplementary. Experiments on diverse 3D scenes demonstrate that TIP-Editor effectively executes various editing tasks, including re-texturing, object insertion, object replacement, and stylization, achieving both high-quality results and strictly following the provided text prompt and reference image.
Keeping unique characteristics specified by the reference image
One of the most distinguishable differences between TIP-Editor and previous methods is that TIP-Editor also supports an image prompt, which offers more accurate control and makes it more user-friendly in real applications. Results in Fig. 1&3 demonstrate high consistency between the updated 3D scene and the reference image (e.g. the styles of the sunglasses; the white giraffe; the virtual ghost horse; the joker make-up appeared in movie The Dark Knight). Moreover, as depicted in the bottom of Fig. 1, our method can also perform global scene editing, such as transferring the entire scene in the Modigliani style of the reference image.
Sequential editing
TIP-Editor can sequentially edit the initial scene multiple times thanks to the local update of the GS and the stepwise 2D personalization strategy, which effectively reduces the interference between the existing scene and the novel content. Results in Fig.4 demonstrate the sequential editing capability. There is no observable quality degradation after multiple times of editing and no interference between different editing operations.
Using generated image as the reference
In the absence of the reference image, we can generate multiple candidates from a T2I model and let the user choose a satisfactory one. This interaction offers the user more control and makes the final result more predictable. Fig. 5 shows some examples.
Figure 6. Visual comparisons between different methods. Our method produces obviously higher-quality results and accurately follows the reference image input (bottom-right corner in column 1). Instruct-N2N sometimes misunderstands (row 1) or overlooks (row 2) the keywords. DreamEditor faces difficulty in making obvious shape changes (row 2). Both of them do not support image prompts to specify detailed appearance/style, producing less controlled results.
Table 1. Quantitative comparisons. CLIPdir denotes the CLIP Text-Image directional similarity. DINOsim is the DINO similarity.
[…]
Instruct-N2N 8.3 36.4 21.6% 8.8%
DreamEditor 11.4 36.8 7.6% 10.0%
Ours 15.5 39.5 70.8% 81.2%
5.3. Comparisons with State-of-the-Art Methods
Qualitative comparisons.
Fig.6 shows visual comparisons between our method and the baselines. Since both baselines do not support image prompts as input, they generate an uncontrolled (probably the most common) item belonging to the object category. In contrast, our results consistently maintain the unique characteristics specified in the reference images (i.e., the heart-shaped sunglasses; the white giraffe; the joker from the movie The Dark Knight).
Moreover, Instruct-N2N sometimes misunderstands (row 1) or overlooks (row 2) the keywords, or cannot generate a specified appearance in limited experiments (row 3), probably due to limited supported instructions in Instruct-Pix2Pix. DreamEditor also faces difficulty if the user wants to add a specified sunglasses item (row 1). Additionally, it is difficult for DreamEditor to make obvious shape changes (row 2) to the existing object due to its adoption of a less flexible mesh-based representation (i.e., NeuMesh).
Quantitative comparisons.
Tab. 1 shows the results of the CLIP Text-Image directional similarity and DINO similarity. The results clearly demonstrate the superiority of our method in both metrics, suggesting that the appearance generated by our method aligns better with both the text prompt and the image prompt. A similar conclusion has been drawn according to the user study. Our results surpass the baselines with a substantial margin on both the quality evaluation ($70.8\%$ votes) and the alignment evaluation ($81.2\%$ votes).
Figure 7. Ablation study on the components proposed in stepwise 2D personalization. We compare the generated images of the personalized T2I model (top row) and the rendered images of the updated 3D scene (bottom row). Removing the localization loss $\mathcal{L}_{loc}$ fails to place the new object in the specified place. Removing the separate LoRA layers dedicated for the personalization of the reference image produces less similar results (heart-shaped vs. regular round shape).
Table 2. Quantitative evaluation of the 3D renderings in the ablation study on the components proposed in stepwise 2D personalization (Fig. 7).
w/o $\mathcal{L}_{loc}$ w/o LoRA Ours
CLIPdir $\uparrow$ 4.4 24.0 25.4
DINOsim $\uparrow$ 18.3 28.0 28.8
Figure 8. Comparison of different 2D personalization methods. Generated images of the T2I models after personalization (top) and the final updated 3D scene (bottom) are presented. Text prompt: “A $V_{1}$ toy wearing $V_{2}$ sunglasses on the forehead”
5.4. Ablation Study
Ablation studies on the stepwise 2D personalization
We conduct ablative experiments in Fig.7 and Tab. 2 to demonstrate the benefit of using $\mathcal{L}_{loc}$ and LoRA Layers in the stepwise 2d personalization. Without $\mathcal{L}_{loc}$, the fine-tuned T2I model fails to place the sunglasses in the specified region (i.e. on the forehead) due to the bias present in the training data of the original T2I model. Introducing dedicated LoRA layers to personalize the unique features in the reference image results in more faithful output, demonstrating the effectiveness of the proposed stepwise 2D personalization strategy in capturing details in the reference image.
We also compare our stepwise 2D personalization with another multiple object personalization method, i.e., Custom Diffusion. As shown in Fig. 8, our stepwise 2D personalization achieves high-quality and faithful personalization results and ensures precise editing within the specified region. In contrast, although Custom Diffusion assigns separate special text tokens to each concept, the UNet is updated by all concepts, resulting in less satisfactory personalization results. In addition, Custom Diffusion lacks a localization mechanism to specify the interaction between the sunglasses and the toy, failing to place the new object on the toy’s forehead.
Figure 9. Ablation study on different 3D representations to show the advantage of GS for this task. Using Instant-NGP results in a changed background while using NeuMesh cannot produce large enough shape deformation. In contrast, using explicit and flexible GS obtains the best foreground editing result while keeping the background unchanged.
Table 3. Quantitative evaluation of the 3D renderings in the ablation study on different 3D representations (Fig. 9).
Instant-NGP NeuMesh Ours
CLIPdir $\uparrow$ 17.8 18.4 18.7
DINOsim $\uparrow$ 31.8 27.0 33.5
Ablation study on different 3D representations
We test different 3D representations in Fig. 9 and Tab. 3 while keeping all the other settings the same. Using GS obtains the best editing result while keeping the background unchanged. For Instant-NGP (Müller et al., 2022), we observe undesired changes in the background since its content in different locations is not independent due to its adoption of a shared MLP and multi-resolution grid. For NeuMesh (Yang et al., 2022) (used in DreamEditor), we observe obvious artifacts if the editing requires (relatively) large shape change since it adopts mesh as its geometry proxy, which is less flexible than GS due to topology constraints.
Figure 10. Ablation study on the influence of global and local SDS (Eq. 8) in the coarse stage. The top row shows the rendering of the editable Gaussians $\mathcal{G^{\mathcal{B}}}$. Only using global SDS $\mathcal{L}^{G}_{SDS}$ produces low-quality foreground object/part, while only using local SDS $\mathcal{L}^{L}_{SDS}$ produces unnatural foreground when composited with the existing scene (e.g., color, placement).
Table 4. Quantitative evaluation of the 3D renderings in the ablation study on different $\gamma$ in coarse editing (Fig. 10).
$\gamma=1$ $\gamma=0.5$ $\gamma=0$
CLIPdir $\uparrow$ 9.0 18.1 12.3
DINOsim $\uparrow$ 29.0 29.5 27.5
Influence of different $\gamma$ in coarse editing
As in Fig.10, both the global and local SDS loss are necessary and our solution achieves the best result. Specifically, only using global SDS loss $\mathcal{L}^{G}_{SDS}$ results in obvious artifacts in the editable region. Only using local SDS loss $\mathcal{L}^{L}_{SDS}$ results in inaccurate placement of the object and unnatural color discrepancy between the background and the novel content since the context information is missing during optimization.
Figure 11. Comparison of the coarse editing results and the refinement results. The region indicated by the arrow demonstrates the efficacy of the refinement step in enhancing the quality of the editing results.
Effectiveness of the pixel-level refinement step
As in Fig.11, introducing the refinement stage effectively reduces artifacts and enhances the texture, resulting in substantially improved quality.
Figure 12. Ablation study on different formulas of $\mathcal{L}_{loc}$. We compare the generated images of the personalized T2I model (top row) and the rendered images of the updated 3D scene (bottom row). Forcing all attention values in the region $\bar{\mathcal{S}}$ to 1 tends to create an object (even multiple objects) covering the entire bounding box regardless of whether its placement is appropriate (e.g. overfill), resulting in various artifacts.
Ablation study on different formulas of $\mathcal{L}_{loc}$
In Eq. 4, we only encourage the max attention value of the pixels in the region $\bar{\mathcal{S}}$ to approach 1. This is different from previous works (Xiao et al., 2023; Avrahami et al., 2023), which force all the attention values in the region $\bar{\mathcal{S}}$ to 1. We compare these two settings in Fig. 12. Using the latter one tends to create an object (even multiple objects) covering the entire bounding box regardless of whether its placement is appropriate (e.g. overfill), resulting in various artifacts.
6. Conclusion and Limitations
In this paper, our proposed TIP-Editor equips the emerging text-driven 3D editing with an additional image prompt as a complement to the textual description and produces high-quality editing results accurately aligned with the text and image prompts while keeping the background unchanged. TIP-Editor offers significantly enhanced controllability and enables versatile applications, including object insertion, object replacement, re-texturing, and stylization.
One limitation of TIP-Editor is the coarse bounding box input. Although convenient, it struggles in complex scenes where bounding boxes may include unwanted elements. It would be very beneficial to automatically obtain 3D instance segmentation of the scene. Another limitation is related to geometry extraction since it is hard to extract a smooth and accurate mesh from GS-represented scenes.
Appendix A More Experimental Results
Figure A.1. The texture details of the rendered image $I_{c}$ from the coarse 3DGS are further enhanced by the denoising process.
Denoising results. Fig. A.1 demonstrates the texture details of the rendered image $I_{c}$ from the coarse 3DGS and its corresponding enhanced image after the denoising process.
Figure A.2. Coarse editing results optimized with different $\gamma$. We demonstrate the rendering images of the edited Gaussian $\mathcal{G^{\mathcal{B}}}$ (top row) and all Gaussians in the scene $\mathcal{G}$ (bottom row). Specifically, only using global SDS loss $\mathcal{L}^{G}_{SDS}$ results in obvious artifacts in the editable region. Coarse editing results optimized with different $\gamma$. We demonstrate the rendering images of the edited Gaussian $\mathcal{G^{\mathcal{B}}}$ (top row) and all Gaussians in the scene $\mathcal{G}$ (bottom row). As the $\gamma$ value increases, global SDS loss $\mathcal{L}^{G}_{SDS}$ obviously results in more artifacts in the editable region. On the contrary, the decreasing of $\gamma$ results in fewer artifacts, but increases the risk of inaccurate object placement and an unnatural color discrepancy between the background and the novel content since the context information is missing during optimization. Generally, setting $\gamma=0.5$ effectively balances $\mathcal{L}^{G}_{SDS}$ and $\mathcal{L}^{L}_{SDS}$, reducing the noise Gaussian while correctly placing the sunglasses.
Influence of the weighting parameter in the coarse stage. Fig. A.2 demonstrates the coarse editing results optimized with different $\gamma$ in Eq. 8. As the $\gamma$ value increases (i.e. increasing the importance of the global SDS loss), more and more obvious artifacts appear in the editable region (i.g. the sunglasses). On the other hand, using too small $\gamma$ values results in fewer artifacts in the sunglasses. However, the final composited image becomes unnatural (e.g. the color gap between the original scene and the new object, the placement of the new object) due to lack of context information during optimization. Generally, setting $\gamma=0.5$ in Eq. 8 strikes a good balances, effectively reducing the noise in the editable Gaussians $\mathcal{G}^{\mathcal{B}}$, correctly placing the sunglasses, and producing harmony colors.
Appendix B Implementation Details
Figure B.1. The generation process of $I_{gt}$ in Sec. 4.3 of the main text. $I_{c}$ denotes the rendered image of the coarse 3DGS. Firstly, we add noise on $I_{c}$ to obtain $I_{c}^{noised}$ and then utilize the personalized T2I model $\epsilon_{\phi_{2}}$ as a denoising network and obtain $I_{c}^{d}$. Secondly, we obtain the binary instance mask $M^{inst}$ of the edited object/part by rendering only the editable Gaussians $\mathcal{G}^{\mathcal{B}}$ and thresholding its opacity mask. Finally, the pseudo GT image $I_{gt}$ is obtained as: $I_{gt}=M^{inst}\odot I_{c}^{d}+(1-M^{inst})\odot I_{bg}$
We provide more implementation details that cannot be included in the main paper due to space. All our experiments are conducted on a single NVIDIA Tesla A100 GPU.
B.1. Optimization of the initial GS scene
In our experiments, we employ 3D Gaussian Splatting (GS) (Kerbl et al., 2023) to represent the 3D scenes. Given images shot from different views of the target scene, we first use COLMAP (Schönberger and Frahm, 2016)) to estimate their corresponding camera parameters. Then, we train the initial GS scene with the default hyper-parameters in the official 3DGS repository (Snosixtyboo, [n. d.]).
B.2. Details of the stepwise 2D personalization strategy
We use the publicly available Stable Diffusion V2.1 (Rombach et al., 2022b) model provided on Hugging Face.
The scene personalization step. We sample scene images with the camera poses covering all elevation and azimuth angles with an interval of $30^{\circ}$ within a predefined range (varied according to the visual scope of the original scene images). We use BLIP-2 (Li et al., 2023) model to obtain the initial scene text prompt, and insert a special token $V_{1}$ in front of the noun to describe this specific scene, resulting in a scene-specific text prompt $y$ during fine-tuning. For the prior preservation loss, we randomly generate 200 images at resolution $512\times 512$ using the same initial scene text $y^{*}$ as input. We use the AdamW optimizer with $lr=5\cdot 10^{-6}$, $beta1=0.9$, $beta2=0.999$, $weightdecay=10^{-2}$, and $epsilon=10^{-8}$ for fine-tuning. The scene personalization step consists of 1K iterations ($\sim 10$ min.) with a batch size of 1. We set $\lambda=0.1$ in $\mathcal{L}_{loc}$ (Eq. 4).
The novel content personalization step. We extract the foreground object in the reference image using an off-the-shelf segmentation model (Ranftl et al., 2021) and resize it into 512$\times$512. We use the AdamW optimizer with $lr=10^{-4}$, $beta1=0.9$, $beta2=0.999$, $weightdecay=10^{-2}$, and $epsilon=10^{-8}$ for fine-tuning. The scene personalization step consists of 500 iterations ($\sim 5$ min.) with a batch size of 1.
B.3. Details of the coarse editing stage
We implement the coarse editing stage in Pytorch (Paszke et al., 2019). We adopt a sampling strategy of views as in (Zhuang et al., 2023), the resolution of the rendered images is 512$\times$512. We use the Adam optimizer with $beta1=0.9$, $beta2=0.999$, $weightdecay=0$, and $epsilon=10^{-15}$ during optimization. We adopt a linear decay strategy for the learning rate optimizing center position $\mu$ from $10^{-3}$ to $2\cdot 10^{-5}$. As for the other attributes, we use a fixed learning rate the same as the official GS settings (i.e. $10^{-1}$ for opacity $\alpha$, $10^{-2}$ for color $c$, and $5\cdot 10^{-2}$ for the 3D covariance matrix $\Sigma$). Since the complexity of different scenes varies significantly, the optimization iterations are varied from 1K ($\sim 5$ min.) to 5K ($\sim 5$ min.) with batch size of 2. For calculating $\mathcal{L}_{SDS}^{G}$ and $\mathcal{L}_{SDS}^{L}$, we set the CFG weight to 10, and utilize simple two-stage annealing of time step $t$: we sample time steps $t\sim\mathcal{U}(0.02,0.75)$ and anneal into $t\sim\mathcal{U}(0.02,0.25)$.
B.4. Details of the refinement stage
Using results from the coarse editing stage as initialization, we continue optimizing the GS with pixel-level image reconstruction loss, which is applied to the rendered image and the carefully generated pseudo GT image $I_{gt}$. Fig. B.1 illustrates the generation of $I_{gt}$. The optimizer and training hyper-parameters are the same as those in training the initial GS scene (Sec. B.1) except for fewer training iterations (3K).
Appendix C Baselines
Instruct-NeRF2NeRF. Instruct-NeRF2NeRF (“I-N2N”) utilizes Instruct-pix2pix (Brooks et al., 2022) to update the rendered multi-view images according to special text instructions obtained with GPT3 (Brown et al., 2020) for supervision. We implement it using its official code.
We use BLIP-2 (Li et al., 2023) to obtain the textual description of our reference image and then convert it into a suitable instruction as in I-N2N. All the other settings are the same as the default configurations in their official code, including the 3D representation method (Nerfacto (Zhang et al., 2021)), the version of Instruct-pix2pix, and the training hyper-parameters.
DreamEditor. DreamEditor adopts the explicit NeuMesh (Yang et al., 2022) representation distilled from a NeuS (Wang et al., 2021) and includes an attention-based localization operation to support local editing. We implement it using its official code and follow its default training hyperparameters. For fair comparisons, we replace its automatic localization with a more accurate manually selected editing area. Moreover, DreamEditor employs Dreambooth to personalize the target scene. We extend its Dreambooth process to simultaneously personalize both the editing scene and the reference image using two different special tokens in one step. We use the same text prompts as our method to edit the scenes.
Appendix D Evaluation Criteria
We compare our method with the aforementioned baselines on 10 results, covering 3 scenes and different editing types. For each result, we render the edited scene for all methods using the same camera trajectory to obtain video demos, which are used for the user study. All the videos are included in this supplementary.
For metric calculation, we extract 1 image every 5 frames from the above videos and calculate the CLIP-based similarity metric (i.e., CLIP Text-Image directional similarity) and DINO similarity. We report their mean values in Tab. 1 in our main paper.
D.1. CLIP Text-Image directional similarity
CLIP Text-Image directional similarity (Gal et al., 2022b) evaluates the degree of alignment between the change of two images (i.e. before-after editing) and two text prompts (i.e. initial text description $t_{o}$ and the editing text prompt $t_{e}$). Mathematically, it is computed as follows:
(D.1) $\begin{split}\Delta T&=E_{T}(t_{e})-E_{T}(t_{o})\\ \Delta I&=E_{I}(i_{e})-E_{I}(i_{o})\\ CLIP_{dir}&=1-\frac{\Delta I\cdot\Delta T}{|\Delta I||\Delta T|}\end{split}$
where $E_{I}$ and $E_{T}$ are CLIP’s image and text encoder, $i_{o}$ and $i_{e}$ are the original and edited scene image, $t_{o}$ and $t_{e}$ are the texts describing the original and edited scene image. The textual descriptions are obtained with BLIP-2 (i.e. $t_{o}$ for $i_{o}$, $t_{r}$ for the reference image). And we manually generate $t_{e}$ according to $t_{o}$ and $t_{r}$. Note that these text descriptions are not necessarily the same as those used for optimization and do not contain special tokens.
D.2. DINO similarity
Following (He et al., 2023), we use DINO (Oquab et al., 2023) similarity measure the similarity between the edited scene and the reference image from different views.
(D.2) $DINO_{sim}=\frac{E_{D}(i_{e})\cdot E_{D}(i_{r})}{|E_{D}(i_{e})||E_{D}(i_{r})|}$
where $E_{D}$ is the image encoder of DINOv2 (Oquab et al., 2023), and $i_{r}$ the Reference image.

### v3

@@
 Extensive experiments have demonstrated that TIP-Editor conducts accurate editing following the text and image prompts in the specified bounding box region, consistently outperforming the baselines in editing quality, and the alignment to the prompts, qualitatively and quantitatively.
-submissionid: 820copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; July 28–August 01, 2024; Denverisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Rendering
+copyright: acmlicensedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 4article: 121publicationmonth: 7doi: 10.1145/3658205ccs: Computing methodologies Rendering
 Figure 1.
@@
 Text-guided image generation and editing
-Text-to-image (T2I) diffusion models  (Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022a), trained on large-scale paired image-text datasets, have gained significant attention since they can generate diverse and high-quality images that match the complicated text prompt.
+Text-to-image (T2I) diffusion models  (Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022), trained on large-scale paired image-text datasets, have gained significant attention since they can generate diverse and high-quality images that match the complicated text prompt.
 Instead of directly generating images from scratch, another popular and closely related task is to edit the given image according to the text prompt (Meng et al., 2021; Couairon et al., 2022; Kawar et al., 2022; Hertz et al., 2022; Avrahami et al., 2022; Brooks et al., 2022).
 Another popular task is object/concept personalization, which aims at generating images for a specified object/concept defined in the given image collection.
-Textual Inversion (TI) (Gal et al., 2022a) optimizes special text token(s) in the text embedding space to represent the specified concept.
+Textual Inversion (TI) (Gal et al., 2022) optimizes special text token(s) in the text embedding space to represent the specified concept.
 DreamBooth (Ruiz et al., 2022) fine-tunes the entire diffusion model with a class-specific prior preservation loss as regularization.
@@
 Recently, text-driven radiance field editing methods have gained more and more attention for their editing flexibility and accessibility.
-For example, Instruct-NeRF2NeRF (Haque et al., 2023) employs an image-based diffusion model (InstructPix2Pix (Brooks et al., 2022)) to modify the rendered image by the users’ instructions, and subsequently update the 3D radiance field with the modified image.
+For example, Instruct-NeRF2NeRF (Haque et al., 2023) and GenN2N (Liu et al., 2024) employ an image-based diffusion model (InstructPix2Pix (Brooks et al., 2022)) to modify the rendered image by the users’ instructions, and subsequently update the 3D radiance field with the modified image.
 DreamEditor (Zhuang et al., 2023), 3D Paintbrush (Decatur et al., 2023) and Vox-E (Sella et al., 2023) enable better local editing by adopting explicit 3D representations (i.e.
@@
 4.1. Stepwise 2D Personalization
-In general, our stepwise personalization of the pre-trained T2I model (i.e., Stable Diffusion (SD) (Rombach et al., 2022a)) is based on DreamBooth (Ruiz et al., 2022), but with two significant modifications.
+In general, our stepwise personalization of the pre-trained T2I model (i.e., Stable Diffusion (SD) (Rombach et al., 2022)) is based on DreamBooth (Ruiz et al., 2022), but with two significant modifications.
 These changes are essential to personalize both the existing scene and the novel content in the reference image.
@@
 1, our method can also perform global scene editing, such as transferring the entire scene in the Modigliani style of the reference image.
+Figure 6.
+Visual comparisons between different methods.
+Our method produces obviously higher-quality results and accurately follows the reference image input (bottom-right corner in column 1).
+Instruct-N2N sometimes misunderstands (row 1) or overlooks (row 2) the keywords.
+DreamEditor faces difficulty in making obvious shape changes (row 2).
+Both of them do not support image prompts to specify detailed appearance/style, producing less controlled results.
 Sequential editing
@@
 5 shows some examples.
-Figure 6.
-Visual comparisons between different methods.
-Our method produces obviously higher-quality results and accurately follows the reference image input (bottom-right corner in column 1).
-Instruct-N2N sometimes misunderstands (row 1) or overlooks (row 2) the keywords.
-DreamEditor faces difficulty in making obvious shape changes (row 2).
-Both of them do not support image prompts to specify detailed appearance/style, producing less controlled results.
 Table 1. Quantitative comparisons. CLIPdir denotes the CLIP Text-Image directional similarity. DINOsim is the DINO similarity.
@@
 Additionally, it is difficult for DreamEditor to make obvious shape changes (row 2) to the existing object due to its adoption of a less flexible mesh-based representation (i.e., NeuMesh).
-Quantitative comparisons.
-Tab.
-1 shows the results of the CLIP Text-Image directional similarity and DINO similarity.
-The results clearly demonstrate the superiority of our method in both metrics, suggesting that the appearance generated by our method aligns better with both the text prompt and the image prompt.
-A similar conclusion has been drawn according to the user study.
-Our results surpass the baselines with a substantial margin on both the quality evaluation ($70.8\%$ votes) and the alignment evaluation ($81.2\%$ votes).
 Figure 7.
@@
 Text prompt: “A $V_{1}$ toy wearing $V_{2}$ sunglasses on the forehead”
+Figure 9.
+Ablation study on different 3D representations to show the advantage of GS for this task.
+Using Instant-NGP results in a changed background while using NeuMesh cannot produce large enough shape deformation.
+In contrast, using explicit and flexible GS obtains the best foreground editing result while keeping the background unchanged.
+Table 3. Quantitative evaluation of the 3D renderings in the ablation study on different 3D representations (Fig. 9).
+Instant-NGP NeuMesh Ours
+CLIPdir $\uparrow$ 17.8 18.4 18.7
+DINOsim $\uparrow$ 31.8 27.0 33.5
+Figure 10.
+Ablation study on the influence of global and local SDS (Eq.
+8) in the coarse stage.
+The top row shows the rendering of the editable Gaussians $\mathcal{G^{\mathcal{B}}}$.
+Only using global SDS $\mathcal{L}^{G}_{SDS}$ produces low-quality foreground object/part, while only using local SDS $\mathcal{L}^{L}_{SDS}$ produces unnatural foreground when composited with the existing scene (e.g., color, placement).
+Table 4. Quantitative evaluation of the 3D renderings in the ablation study on different $\gamma$ in coarse editing (Fig. 10).
+$\gamma=1$ $\gamma=0.5$ $\gamma=0$
+CLIPdir $\uparrow$ 9.0 18.1 12.3
+DINOsim $\uparrow$ 29.0 29.5 27.5
+Quantitative comparisons.
+Tab.
+1 shows the results of the CLIP Text-Image directional similarity and DINO similarity.
+The results clearly demonstrate the superiority of our method in both metrics, suggesting that the appearance generated by our method aligns better with both the text prompt and the image prompt.
+A similar conclusion has been drawn according to the user study.
+Our results surpass the baselines with a substantial margin on both the quality evaluation ($70.8\%$ votes) and the alignment evaluation ($81.2\%$ votes).
 5.4. Ablation Study
@@
 In addition, Custom Diffusion lacks a localization mechanism to specify the interaction between the sunglasses and the toy, failing to place the new object on the toy’s forehead.
-Figure 9.
-Ablation study on different 3D representations to show the advantage of GS for this task.
-Using Instant-NGP results in a changed background while using NeuMesh cannot produce large enough shape deformation.
-In contrast, using explicit and flexible GS obtains the best foreground editing result while keeping the background unchanged.
-Table 3. Quantitative evaluation of the 3D renderings in the ablation study on different 3D representations (Fig. 9).
-Instant-NGP NeuMesh Ours
-CLIPdir $\uparrow$ 17.8 18.4 18.7
-DINOsim $\uparrow$ 31.8 27.0 33.5
 Ablation study on different 3D representations
@@
 Figure 11.
-Ablation study on the influence of global and local SDS (Eq.
-8) in the coarse stage.
-The top row shows the rendering of the editable Gaussians $\mathcal{G^{\mathcal{B}}}$.
-Only using global SDS $\mathcal{L}^{G}_{SDS}$ produces low-quality foreground object/part, while only using local SDS $\mathcal{L}^{L}_{SDS}$ produces unnatural foreground when composited with the existing scene (e.g., color, placement).
-Table 4. Quantitative evaluation of the 3D renderings in the ablation study on different $\gamma$ in coarse editing (Fig. 10).
-$\gamma=1$ $\gamma=0.5$ $\gamma=0$
-CLIPdir $\uparrow$ 9.0 18.1 12.3
-DINOsim $\uparrow$ 29.0 29.5 27.5
+Comparison of the coarse editing results and the refinement results.
+The region indicated by the arrow demonstrates the efficacy of the refinement step in enhancing the quality of the editing results.
 Influence of different $\gamma$ in coarse editing
@@
 Only using local SDS loss $\mathcal{L}^{L}_{SDS}$ results in inaccurate placement of the object and unnatural color discrepancy between the background and the novel content since the context information is missing during optimization.
-Figure 11.
-Comparison of the coarse editing results and the refinement results.
-The region indicated by the arrow demonstrates the efficacy of the refinement step in enhancing the quality of the editing results.
 Effectiveness of the pixel-level refinement step
@@
 Another limitation is related to geometry extraction since it is hard to extract a smooth and accurate mesh from GS-represented scenes.
-Appendix A More Experimental Results
-Figure A.1. The texture details of the rendered image $I_{c}$ from the coarse 3DGS are further enhanced by the denoising process.
-Denoising results. Fig. A.1 demonstrates the texture details of the rendered image $I_{c}$ from the coarse 3DGS and its corresponding enhanced image after the denoising process.
-Figure A.2.
-Coarse editing results optimized with different $\gamma$.
-We demonstrate the rendering images of the edited Gaussian $\mathcal{G^{\mathcal{B}}}$ (top row) and all Gaussians in the scene $\mathcal{G}$ (bottom row).
-Specifically, only using global SDS loss $\mathcal{L}^{G}_{SDS}$ results in obvious artifacts in the editable region.
-Coarse editing results optimized with different $\gamma$.
-We demonstrate the rendering images of the edited Gaussian $\mathcal{G^{\mathcal{B}}}$ (top row) and all Gaussians in the scene $\mathcal{G}$ (bottom row).
-As the $\gamma$ value increases, global SDS loss $\mathcal{L}^{G}_{SDS}$ obviously results in more artifacts in the editable region.
-On the contrary, the decreasing of $\gamma$ results in fewer artifacts, but increases the risk of inaccurate object placement and an unnatural color discrepancy between the background and the novel content since the context information is missing during optimization.
-Generally, setting $\gamma=0.5$ effectively balances $\mathcal{L}^{G}_{SDS}$ and $\mathcal{L}^{L}_{SDS}$, reducing the noise Gaussian while correctly placing the sunglasses.
-Influence of the weighting parameter in the coarse stage.
-Fig.
-A.2 demonstrates the coarse editing results optimized with different $\gamma$ in Eq.
-8.
-As the $\gamma$ value increases (i.e.
-increasing the importance of the global SDS loss), more and more obvious artifacts appear in the editable region (i.g.
-the sunglasses).
-On the other hand, using too small $\gamma$ values results in fewer artifacts in the sunglasses.
-However, the final composited image becomes unnatural (e.g.
-the color gap between the original scene and the new object, the placement of the new object) due to lack of context information during optimization.
-Generally, setting $\gamma=0.5$ in Eq.
-8 strikes a good balances, effectively reducing the noise in the editable Gaussians $\mathcal{G}^{\mathcal{B}}$, correctly placing the sunglasses, and producing harmony colors.
-Appendix B Implementation Details
-Figure B.1.
-The generation process of $I_{gt}$ in Sec.
-4.3 of the main text.
-$I_{c}$ denotes the rendered image of the coarse 3DGS.
-Firstly, we add noise on $I_{c}$ to obtain $I_{c}^{noised}$ and then utilize the personalized T2I model $\epsilon_{\phi_{2}}$ as a denoising network and obtain $I_{c}^{d}$.
-Secondly, we obtain the binary instance mask $M^{inst}$ of the edited object/part by rendering only the editable Gaussians $\mathcal{G}^{\mathcal{B}}$ and thresholding its opacity mask.
-Finally, the pseudo GT image $I_{gt}$ is obtained as: $I_{gt}=M^{inst}\odot I_{c}^{d}+(1-M^{inst})\odot I_{bg}$
-We provide more implementation details that cannot be included in the main paper due to space. All our experiments are conducted on a single NVIDIA Tesla A100 GPU.
-B.1. Optimization of the initial GS scene
-In our experiments, we employ 3D Gaussian Splatting (GS) (Kerbl et al., 2023) to represent the 3D scenes.
-Given images shot from different views of the target scene, we first use COLMAP (Schönberger and Frahm, 2016)) to estimate their corresponding camera parameters.
-Then, we train the initial GS scene with the default hyper-parameters in the official 3DGS repository (Snosixtyboo, [n.
-d.]).
-B.2. Details of the stepwise 2D personalization strategy
-We use the publicly available Stable Diffusion V2.1 (Rombach et al., 2022b) model provided on Hugging Face.
-The scene personalization step.
-We sample scene images with the camera poses covering all elevation and azimuth angles with an interval of $30^{\circ}$ within a predefined range (varied according to the visual scope of the original scene images).
-We use BLIP-2 (Li et al., 2023) model to obtain the initial scene text prompt, and insert a special token $V_{1}$ in front of the noun to describe this specific scene, resulting in a scene-specific text prompt $y$ during fine-tuning.
-For the prior preservation loss, we randomly generate 200 images at resolution $512\times 512$ using the same initial scene text $y^{*}$ as input.
-We use the AdamW optimizer with $lr=5\cdot 10^{-6}$, $beta1=0.9$, $beta2=0.999$, $weightdecay=10^{-2}$, and $epsilon=10^{-8}$ for fine-tuning.
-The scene personalization step consists of 1K iterations ($\sim 10$ min.) with a batch size of 1.
-We set $\lambda=0.1$ in $\mathcal{L}_{loc}$ (Eq.
-4).
-The novel content personalization step.
-We extract the foreground object in the reference image using an off-the-shelf segmentation model (Ranftl et al., 2021) and resize it into 512$\times$512.
-We use the AdamW optimizer with $lr=10^{-4}$, $beta1=0.9$, $beta2=0.999$, $weightdecay=10^{-2}$, and $epsilon=10^{-8}$ for fine-tuning.
-The scene personalization step consists of 500 iterations ($\sim 5$ min.) with a batch size of 1.
-B.3. Details of the coarse editing stage
-We implement the coarse editing stage in Pytorch (Paszke et al., 2019).
-We adopt a sampling strategy of views as in (Zhuang et al., 2023), the resolution of the rendered images is 512$\times$512.
-We use the Adam optimizer with $beta1=0.9$, $beta2=0.999$, $weightdecay=0$, and $epsilon=10^{-15}$ during optimization.
-We adopt a linear decay strategy for the learning rate optimizing center position $\mu$ from $10^{-3}$ to $2\cdot 10^{-5}$.
-As for the other attributes, we use a fixed learning rate the same as the official GS settings (i.e.
-$10^{-1}$ for opacity $\alpha$, $10^{-2}$ for color $c$, and $5\cdot 10^{-2}$ for the 3D covariance matrix $\Sigma$).
-Since the complexity of different scenes varies significantly, the optimization iterations are varied from 1K ($\sim 5$ min.) to 5K ($\sim 5$ min.) with batch size of 2.
-For calculating $\mathcal{L}_{SDS}^{G}$ and $\mathcal{L}_{SDS}^{L}$, we set the CFG weight to 10, and utilize simple two-stage annealing of time step $t$: we sample time steps $t\sim\mathcal{U}(0.02,0.75)$ and anneal into $t\sim\mathcal{U}(0.02,0.25)$.
-B.4. Details of the refinement stage
-Using results from the coarse editing stage as initialization, we continue optimizing the GS with pixel-level image reconstruction loss, which is applied to the rendered image and the carefully generated pseudo GT image $I_{gt}$.
-Fig.
-B.1 illustrates the generation of $I_{gt}$.
-The optimizer and training hyper-parameters are the same as those in training the initial GS scene (Sec.
-B.1) except for fewer training iterations (3K).
-Appendix C Baselines
-Instruct-NeRF2NeRF.
-Instruct-NeRF2NeRF (“I-N2N”) utilizes Instruct-pix2pix (Brooks et al., 2022) to update the rendered multi-view images according to special text instructions obtained with GPT3 (Brown et al., 2020) for supervision.
-We implement it using its official code.
-We use BLIP-2 (Li et al., 2023) to obtain the textual description of our reference image and then convert it into a suitable instruction as in I-N2N.
-All the other settings are the same as the default configurations in their official code, including the 3D representation method (Nerfacto (Zhang et al., 2021)), the version of Instruct-pix2pix, and the training hyper-parameters.
-DreamEditor.
-DreamEditor adopts the explicit NeuMesh (Yang et al., 2022) representation distilled from a NeuS (Wang et al., 2021) and includes an attention-based localization operation to support local editing.
-We implement it using its official code and follow its default training hyperparameters.
-For fair comparisons, we replace its automatic localization with a more accurate manually selected editing area.
-Moreover, DreamEditor employs Dreambooth to personalize the target scene.
-We extend its Dreambooth process to simultaneously personalize both the editing scene and the reference image using two different special tokens in one step.
-We use the same text prompts as our method to edit the scenes.
-Appendix D Evaluation Criteria
-We compare our method with the aforementioned baselines on 10 results, covering 3 scenes and different editing types.
-For each result, we render the edited scene for all methods using the same camera trajectory to obtain video demos, which are used for the user study.
-All the videos are included in this supplementary.
-For metric calculation, we extract 1 image every 5 frames from the above videos and calculate the CLIP-based similarity metric (i.e., CLIP Text-Image directional similarity) and DINO similarity.
-We report their mean values in Tab.
-1 in our main paper.
-D.1. CLIP Text-Image directional similarity
-CLIP Text-Image directional similarity (Gal et al., 2022b) evaluates the degree of alignment between the change of two images (i.e.
-before-after editing) and two text prompts (i.e.
-initial text description $t_{o}$ and the editing text prompt $t_{e}$).
-Mathematically, it is computed as follows:
-(D.1) $\begin{split}\Delta T&=E_{T}(t_{e})-E_{T}(t_{o})\\ \Delta I&=E_{I}(i_{e})-E_{I}(i_{o})\\ CLIP_{dir}&=1-\frac{\Delta I\cdot\Delta T}{|\Delta I||\Delta T|}\end{split}$
-where $E_{I}$ and $E_{T}$ are CLIP’s image and text encoder, $i_{o}$ and $i_{e}$ are the original and edited scene image, $t_{o}$ and $t_{e}$ are the texts describing the original and edited scene image.
-The textual descriptions are obtained with BLIP-2 (i.e.
-$t_{o}$ for $i_{o}$, $t_{r}$ for the reference image).
-And we manually generate $t_{e}$ according to $t_{o}$ and $t_{r}$.
-Note that these text descriptions are not necessarily the same as those used for optimization and do not contain special tokens.
-D.2. DINO similarity
-Following (He et al., 2023), we use DINO (Oquab et al., 2023) similarity measure the similarity between the edited scene and the reference image from different views.
-(D.2) $DINO_{sim}=\frac{E_{D}(i_{e})\cdot E_{D}(i_{r})}{|E_{D}(i_{e})||E_{D}(i_{r})|}$
-where $E_{D}$ is the image encoder of DINOv2 (Oquab et al., 2023), and $i_{r}$ the Reference image.
+Acknowledgements.
+This work was supported in part by the National Natural Science Foundation of China (NO.
+62325605, NO.
+62322608), in part by the CAAI-MindSpore Open Fund, developed on OpenI Community,in part by the Open Project Program of State Key Laboratory of Virtual Reality Technology and Systems, Beihang University (No.VRLAB2023A01).